Flawless Cosmetic Company manufactures and distributes several different products. The company currently uses a plantwide allocation method for allocating overhead at a rate of $5 per direct labor hour. Loren is the department manager of the Makeup Department which produces Products – Concealer (C) and Glow Cream (GC). Jennifer is the department manager of the Hair Care Department which manufactures Product – Shampoo (S). The product costs (per case of 24 bottles) and other information are as follows:

Products
C GC S
Direct materials $ 120.00 $ 72.00 $ 46.00
Direct labor 50.00 34.00 12.00
Overhead 34.00 18.00 14.00
$ 204.00 $ 124.00 $ 72.00
Machine hours 4 2 3
Number of cases (per year) 300 500 600

Based on this information, if Flawless changes its allocation basis to machine hours, what is the overhead rate per machine hour?

$2.00

$3.00

$6.900

$18.00
Answer #1

Total Overhead

= (34*300)+(18*500)+(14*600)

= 27,600

Total machine hours

= (300*4)+(500*2)+(600*3)

= 4,000

Overhead rate = Total Overhead/Total machine hours

= 27,600/4000

= 6.900